149
edits
No edit summary |
No edit summary |
||
--[[User:Zjsimmons|Zjsimmons]] ([[User talk:Zjsimmons|talk]]) 18:33, 14 July 2014 (CDT)--[[User:Zjsimmons|Zjsimmons]] ([[User talk:Zjsimmons|talk]]) 18:33, 14 July 2014 (CDT)
Made some changes today, see qualitative improvement in green lock. So we've been having issues with the locking performance, the transmitted IR and also generated green were intensity modulated at about 300Hz, a strange freq to see problems as it was not the cavity piezo resonance freq of about 420Hz. The SHG derived green looked at times almost like a pulse train as a result of the modulation. This was tracked to in part a suspected ground loop between driver and locking electronics. Unfortunately the fast feedback was needed to get good output power, but would also cause the modulation.
*Swapped the cavity piezo for the new shorter travel (6um) one from Noliac. Glued everything together with epoxy so as to cut down on load weight. Somehow, realigning the cavity didn't turn into a huge project. Note: when you see some green, put the spots on top of each other and you're most of the way toward good alignment.
*Performed the resonance test on the piezo/mirror system. We now see resonance at about 700Hz. This is higher but not dramatically so. Also, the resonance was qualitatively not as noticeable, so perhaps it is less strong as well which would be cool.
Upon locking with the new piezo, observed very robust stable lock, you can really rap on the table and it won't drop. Piezo/mirror system definitely less susceptible to noise, banging, disturbing the lock. However we still see the modulation present via the fast feedback. Turning the FFB on and tuning to the best gain the output looks pretty good. Would estimate fluctuation in about the top 25% of the transmitted IR or so. Turning the piezo gain all the way up saw no piezo resonance with FFB off. Green is relatively constant, i.e. not intermittent, although there is fluctuation in the top 1/2 or 1/3 so.
*Swapped R30 in the circuit from 1k to 100k, this should boost the piezo feedback by a factor of 100. Since we saw no resonance at max piezo gain with the new piezo, i decided to change the resistor to give us more gain to play with. Now at higher piezo gain (adjusted via pot) we can observe piezo resonance. FFB squelches this but we still see some wiggling on the transmitted IR and Green. Now the wiggles are more like 100-150Hz. There also seems to be a tad more fluctuation in general at low piezo gain now. I split the difference and change the 100k to 30k.
*Next priority: find some way to squelch the FFB noise problem. Looking online, and from Deniz's comments, it seems that the laser driver mode input is very susceptible to noise/ground loop issues.
--[[User:Zjsimmons|Zjsimmons]] ([[User talk:Zjsimmons|talk]]) 21:54, 11 July 2014 (CDT)
Week in review: Frustratingly slow week; didn't seem to make much headway/i don't know anything.
|